OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? point 8,-4 and slope -5/6

OpenStudy (perl):

looks correct

OpenStudy (anonymous):

sorry link doesn't work

OpenStudy (anonymous):

That's correct.

OpenStudy (anonymous):

HERE'S MORE; A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks: 07_08_LQ_4.gif Which equation gives the cost, y, in terms of the number of pages, x? A. y = 0.012x + 49.985 B. y = 0.012x + 1.25 C. y = 83.3x + 0.65 D. y = 0.012x + 0.65

OpenStudy (anonymous):

tHERE'S THE LINK

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@iRespect